# HHardware specifications

### v1

@@ 3.2.3 Continuous Variable QKD
The study performed in this section pertains to CV-QKD protocols, but could be adapted to coherent classical communication protocols, as they use similar hardware. Indeed, the usual setup is based on standard telecom technologies, i.e., at a wavelength of $1550$ nm, and Balanced Homodyne Detectors (BHDs) with typical efficiencies above $p_{det}=0.7$, and electronic noise around $V_{el}=0.005$ SNU, in agreement with recent advances in the field [29, 30]. Other parameters used for this section can be found in Table 1. In particular, we choose a higher source rate $r_{\rm source}=100$ MHz than in the case of DV-protocol, which is representative of the latest CV-QKD experiments [23].
@@
Figure 5(a) shows $E^{\rm 1Gbit}_{\text{Gaussian CV-QKD}}$ for homodyne- and heterodyne-based detections, as well as single- and double-polarizations. It can be observed that, for short and medium distances, the overall energy consumption remains constant and is given by the startup energy of the setup. For distances beyond 100 km, however, there is a clear advantage in using a double-polarization scheme since it reduces the overall execution time of the protocol without noticeably increasing the power consumption. Furthermore, the homodyne approach provides a slight improvement in terms of energy for very long distances. The similarity between the homodyne and heterodyne cases is explained by the fact that, while a heterodyne detection measures both quadratures for every round which increases the rate, it then requires two homodyne detection setups which increases the energy cost. Said cost could be reduced by using the RF heterodyne detection scheme, where the two quadratures can be measured with a single balanced detector [31]. These two effects were observed to mostly cancel out, with a slight outperformance of the heterodyne setup. This could nonetheless represent an advantage in terms of scalability for networks of many users.
@@
(a) Energy required to distill 1 Gbit of secret key using a Gaussian modulation. The variance of the states was optimized with respect to the distance.
(b) Energy required to distill 1 Gbit of secret key for the QPSK scheme. The amplitude of the states was optimized with respect to the distance.
Figure 5: Study of the energetic consumption of CV-QKD protocols.
@@
The energy consumption of the QPSK protocol is shown in Figure 5(b). The experimental setup is identical to the Gaussian modulation scheme with the corresponding detection apparatus. As in the previous case, the double-polarization provides a clear reduction in the energy consumption, and both homodyne and heterodyne detections provide an overall similar performance. At low distances, the energy consumption does not differ compared to using Gaussian-modulated states. This is explained by the fact that the same hardware setup is considered for both Gaussian and QPSK modulation, and the latter has a considerably lower key rate at high distance. This advantage of the Gaussian distribution is however affected by the inclusion of the cost of classical post-processing, as the digital signal processing and error reconciliation steps may prove to be less energy costly for discrete modulations. Other discrete modulations may also prove more efficient such as Quadrature-Amplitude Modulation, which may make use of a Probabilistic Constellation Shaping [32].
3.3 Classical costs and comparison between CV-QKD and DV-QKD
@@
Results from Figure 4(a) and Figure 5(a) hint at CV-QKD being less energy consuming than any DV-QKD protocols. However, in order to get a meaningful comparison between DV- and CV-QKD protocols, one has also to consider the energetic costs of classical post-processing, which are referred to as classical costs in the rest of this section. In addition to being challenging to estimate, they also largely differ from one family of protocols to the other.
@@
In practice, the biggest difference between CV- and DV-QKD is the digital signal processing (DSP). Indeed, in the DV case, where the photons are detected using single photon detectors, the signal processing is mostly done by the time tagger, that has a known energy consumption and is taken into account in the DV setups. On the other side, recent CV-QKD setups ([23, 31, 33]) are using advanced digital signal processing techniques, which are more costly in energy, and cannot be fully realized in real-time at the time of writing of this paper. Note that information reconciliation cost may also differ significantly between CV- and DV-QKD, as error correction for complex variables is more involved than the one for binary variables. However, estimating the difference between the two costs is not trivial, and is not considered in this analysis.
In CV-QKD, the signal is acquired by the Analog-to-Digital Converter (ADC), and then a series of filters and classical algorithms are applied to recover the symbols. We make the assumption that the cost to recover one symbol from the original signal is a constant and denote it as $\tau_{DSP}$. The energy contribution from signal processing is then given by $\tau_{DSP}$ multiplied by the number of symbols exchanged over the quantum channels:
@@
where $N_{\mathrm{target}}$ is the target number of bits in the final secret key and $K_{\mathrm{CV-QKD}}$ is the secret key rate of the CV-QKD protocol (in bit/symbol).
To get an estimate of $\tau_{DSP}$, the open source QOSST software [31] is used as reference, where the DSP runs on a computer during 3 min for 1 million symbols. Assuming a power of 100 W for the computer, this gives an already-achieved value for $\tau_{DSP}=0.018$ J/symbol. Since the software is written in Python, its running time could be optimized greatly, and a value of 1 min for 1 million symbols could be reached in the near future, which would give a slightly more optimistic value $\tau_{DSP}=0.006$ J/symbol.
The costs of the BB84 protocol with APDs and with SNSPDs, as well as the costs of Gaussian-modulated CV-QKD without DSP and for $\tau_{DSP}=0.006$ or $0.018$ J/symbol are shown in Figure 6 . The results reveal that, when neglecting the DSP cost, CV-QKD always outperform BB84. However, with the DSP cost, there is only a small regime of distances (below 5 km) where CV-QKD is more efficient than BB84 with APDs. For high distances ($>75$ km) even BB84 with SNSPDs consumes less than CV-QKD.
@@
4 Energy cost of multipartite protocols
Protocols involving three or more nodes in a quantum network are of particular interest for future deployments of the quantum internet. Not all quantum technologies scale linearly, and therefore it is worth finding, energy-wise, optimal configurations for multipartite networks. Firstly, different methods of generation of all-to-all entanglement are described and compared. Secondly, Conference Key Agreement (CKA) is used as a figure of merit for multipartite networks by comparing the performances of a few protocols constructed from DV and CV sources.
4.1 All-to-all entanglement
@@
GHZ states are prime candidates for network applications since they allow the sharing of entanglement between all nodes at once, as seen in many network protocols [10, 12, 36]. A common photonic GHZ state creation setup involves SPDC sources creating Bell pairs that go through a fusion operation which is inherently probabilistic (see Appendix B.3). As a consequence, the probability of successfully creating a GHZ state of $n$ qubits decreases exponentially with $n$.
@@
$E_{\rm GHZ}(n,t)=\left\lceil{\frac{n}{2}}\right\rceil\;S_{\textrm{SPDC}}(t)+\left\lfloor{\frac{n-1}{2}}\right\rfloor\;M_{\rm fusion,polarization}(t)+n\;D_{\rm SNSPD}(t)+C_{\rm GHZ}(t),$ (7)
@@
$E_{\rm Alltoall}(n,t)=\frac{n(n-1)}{2}\;S_{\textrm{SPDC}}(t)+n(n-1)\;M_{\rm polar}(t)+n\;D_{\rm SNSPD}(t)+C_{\rm Alltoall}(t).$ (8)
While this second method amounts to more hardware involved, the low probability of success of high order GHZ state creation requires more repetitions and thus longer running times of the hardware. As shown in Figure 7, the best method to share entanglement between all nodes of a network varies with the number of parties. After $n=6$ parties, the probability of successfully creating a GHZ state becomes so low that it is better, energy wise, to use a pair-wise entangled architecture for this task.
@@
4.2 Conference key agreement
Conference key agreement (CKA) [37] is the multipartite extension of QKD, allowing $n$ parties to create and share a common secret key. Similarly to QKD, there are several protocols that achieve CKA, each with different pros and cons in terms of rate, security bound, and, as per the purpose of this work, energy consumption.
@@ 4.2.1 DV-CKA
The most straightforward method to create a secret key between $n$ users of a quantum network is to distribute the qubits of a GHZ state to each party of a network. By measuring their qubits in the appropriate bases, the parties automatically extract a common bit, whose privacy follows from the monogamy of entanglement. A full security proof and description of a GHZ-based CKA protocol can be found in [38], while an experimental realization can be found in [39]. The rate of this protocol, which we denote GHZ-CKA in the following section, is given by the rate of creating, sharing, and measuring GHZ states that were distributed to the $n$ parties. Assuming that all $n$ parties are at an equal distance $d$ from the source of GHZ states, Equation 4 becomes:
@@ 4.2.2 CV-CKA
A brief description of the method to compute the secret key rate of the protocol [40] is given in Appendix C.2.3. As a concluding remark, each of the Bobs needs to perform post-processing of the measured outcomes, which means that the classical cost increases with the number of users (see Section 3.3).
@@ 4.2.3 Simulation results for CKA
(a) Energy consumption of the three different CKA setups as a function of the number of parties for a fixed distance between the parties and the central node of 5 km, in log-scale.
(b) Energy consumption of the three different CKA setups as a function of the distance between the parties and the central node, for a fixed number of parties $n=5$, in log-scale.
Figure 8: Energetic analysis of DV-CKA protocols.
In Figures 8(a) and 8(b), the energy cost $E^{\textrm{ \rm 1Gbit}}_{\textrm{CKA}}$ required to create 1 Gbit of key is illustrated, using the three aforementioned DV-CKA protocols as a function of the number of parties and as a function of the distance with the middle party Alice. The BB84-CKA protocol is always more energy efficient than the other two options. This is due to the high rate and high success probability of the BB84 protocol. As for the All-to-all task, there is a regime for both the number of parties and for the distance for which the GHZ-CKA protocol is more energy efficient than the Bell-CKA protocol. For longer distances and for higher numbers of parties, the probability of a successful generation and detection of a GHZ state decreases to the point where it becomes more efficient to use bipartite communications between Alice and each Bob to accomplish this task.
@@
(a) Energy consumption required to distill 1Gbit of key with the CV-CKA protocol as a function of the distance with the central node, for different numbers of users.
(b) Energy consumption required to distill 1Gbit of key with the nCV-QKD protocol (using homodyne measurements and double-polarization) as a function of the distance with the central node, for different numbers of users.
Figure 9: Energetic analysis of CV-CKA protocols.
The energy consumption required to create 1 Gbit of secret key using the CV-CKA protocol of [40] as a function of distance is shown in Figure 9(a) for different numbers of parties. Regarding the comparison between continuous and discrete variables, the energy consumption of the CV-CKA protocol is three orders of magnitude lower than its DV-counterpart, but with major limitations on the achievable distance since this energetic advantage is valid only over a few hundred meters [40]. This is therefore a solution to be considered for a few users within building-scale distances, compared to other DV-CKA protocols presented in previous sections.
Figure 9(b) shows the energy consumption required to create 1 Gbit of secret key using the nCV-QKD protocol for diverse numbers of users $n$, all of them separated by the same distance. A similar energy consumption with respect to the CV-CKA approach is observed with a noticeable improvement in terms of achievable distances. Both techniques employ vastly different classical post-processing (in particular, the DSP), such that adding those contributions could modify the results of this simulation.
5 Conclusion
In this article, we laid the foundations of a framework to estimate the energy consumption of quantum communication protocols. Two figures of merit were introduced, namely the energy efficiency and the energy cost required to produce a target number of secret bits. We applied them to different implementations of bipartite and multipartite protocols. The energy efficiency gives an idea of the consumption of a running network while the energy cost of producing a target number of bits gives a benchmark for the current and near-term energy cost of network protocols. By studying the energy cost required to solve identical basic tasks through different protocols and hardware choices, hardware and protocol choices can be optimized.
This first insight into the energetic cost of photonic-based quantum communication protocols, including multipartite scenarios, shows that the critical components in discrete variable protocols are the cryogenic-based hardware, while continuous variable protocols are deeply affected by post-processing. An interesting highlight of this study is that, for a distance of 25 km, a typical BB84 setup could generate 1 Gbit of secret key 60 times more efficiently using a less energy costly but less efficient detector than the usual cryogenic-based detectors, at the cost of 3 times the temporal requirements. These differences are critical at this early stage of quantum network development and can direct future efforts in different directions.

### v2

@@ III.2.3 Continuous Variable QKD
The study in this section pertains to CV-QKD protocols, but could be adapted to coherent classical communication protocols, as they use similar hardware. Indeed, the usual setup is based on standard telecom technologies, i.e., at a wavelength of $1550$ nm, and Balanced Homodyne Detectors (BHDs) with typical efficiencies above $p_{det}=0.7$, and electronic noise around $V_{el}=0.005$ SNU, in agreement with recent advances in the field [28, 29]. Other parameters used for this section can be found in Table 1. In particular, we choose a higher source rate $r_{\rm source}=100$ MHz than in the case of DV-protocol, which is representative of the latest CV-QKD experiments [22].
@@
Figure 6 shows $E^{\rm 1Gbit}_{\text{Gaussian CV-QKD}}$ for homodyne- and heterodyne-based detections, as well as single- and double-polarizations. It can be observed that, for short and medium distances, the overall energy consumption remains constant and is given by the startup energy of the setup. For distances beyond 100 km, however, there is a clear advantage in using a double-polarization scheme since it reduces the overall execution time of the protocol without noticeably increasing the power consumption. Furthermore, the homodyne approach provides a slight improvement in terms of energy for very long distances. The similarity between the homodyne and heterodyne cases is explained by the fact that, while a heterodyne detection measures both quadratures for every round which increases the rate, it then requires two homodyne detection setups which increases the energy cost. Said cost could be reduced by using the RF heterodyne detection scheme, where the two quadratures can be measured with a single balanced detector [30]. These two effects were observed to mostly cancel out, with a slight outperformance of the heterodyne setup. This could nonetheless represent an advantage in terms of scalability for networks of many users.
[…]
@@
Figure 6: Study of the energetic consumption of CV-QKD protocols. (a) Energy required to distill 1 Gbit of secret key using a Gaussian modulation. (b) Energy required to distill 1 Gbit of secret key for the QPSK scheme. In both cases, the variance of the states was optimized with respect to the distance. In both plots, the orange line is superposed with the red one and the blue line is superposed with the green one.
@@
The energy consumption of the QPSK protocol is shown in Figure 6. The experimental setup is identical to the Gaussian modulation scheme with the corresponding detection apparatus. As in the previous case, the double-polarization provides a clear reduction in the energy consumption, and both homodyne and heterodyne detections provide an overall similar performance. At low distances, the energy consumption does not differ compared to using Gaussian-modulated states. This is explained by the fact that the same hardware setup is considered for both Gaussian and QPSK modulation, and the latter has a considerably lower key rate at high distance. This advantage of the Gaussian distribution is however affected by the inclusion of the cost of classical post-processing, as the digital signal processing and error reconciliation steps may prove to be less energy costly for discrete modulations. Other discrete modulations may also prove more efficient such as Quadrature-Amplitude Modulation, which may make use of a Probabilistic Constellation Shaping [31].
III.3 Classical costs and comparison between CV-QKD and DV-QKD
@@
Results from Figure 4 and Figure 6 hint at CV-QKD being less energy consuming than any DV-QKD protocols. However, in order to get a meaningful comparison between DV- and CV-QKD protocols, one has also to consider the energetic costs of classical post-processing, which are referred to as classical costs in the rest of this section. In addition to being challenging to estimate, they also largely differ from one family of protocols to the other.
@@
In practice, the biggest difference between CV- and DV-QKD is the digital signal processing (DSP). Indeed, in the DV case, where the photons are detected using single photon detectors, the signal processing is partly done by the time tagger, that has a known energy consumption and is taken into account in the DV setups. On the other side, recent CV-QKD setups ([22, 30, 32]) are using advanced digital signal processing techniques, which are more costly in energy and cannot be fully realized in real-time at the time of writing of this paper. In this study, we take into account this DSP cost for CV-QKD. Note, however, that the DSP handles a certain number of impairment corrections, also accounting for synchronization (in time, frequency, phase, and sometimes in polarization) which is also needed in DV-QKD experiments. In this sense, we are neglecting some of the DV classical costs, which implies a slight underestimation compared to the cost of CV-QKD. Note also that the information reconciliation cost may differ significantly between CV- and DV-QKD, as error correction for complex variables is more involved than for binary variables. However, estimating the difference between the two costs is not trivial, and is not considered in this analysis.
In CV-QKD, the signal is acquired by the Analog-to-Digital Converter (ADC), and then a series of filters and classical algorithms are applied to recover the symbols. We assume that the cost to recover one symbol from the original signal is a constant and denote it as $\tau_{DSP}$. The energy contribution from signal processing is then given by $\tau_{DSP}$ multiplied by the number of symbols exchanged over the quantum channels:
@@
where $N_{\mathrm{target}}$ is the target number of bits in the final secret key and $K_{\mathrm{CV-QKD}}$ is the secret key rate of the CV-QKD protocol (in bit/symbol). This model was validated on the experimental setup of [30] but might not be the most general one.
@@
To get an estimate of $\tau_{DSP}$, the open source QOSST software [30] is used as reference, where the DSP runs on a computer during 3 min for 1 million symbols. Assuming a power of 100 W for the computer, this gives an already-achieved value for $\tau_{DSP}=0.018$ J/symbol. Since the software is written in Python, its running time could be optimized greatly, and a value of 1 min for 1 million symbols could be reached in the near future, which would give a slightly more optimistic value $\tau_{DSP}=0.006$ J/symbol. Note that since the publication of [30], improvements on the DSP algorithm done partially under the impulsion of the preliminary results of this study allowed the DSP to run in 3 seconds for 1 million symbols on average, giving a $\tau_{DSP}=0.0003$ J/symbol.
The costs of the BB84 protocol with APDs and with SNSPDs, as well as the costs of Gaussian-modulated CV-QKD without DSP and for the different values of $\tau_{DSP}$ are shown in Figure 7. The results reveal that, when neglecting the DSP cost, CV-QKD always outperform BB84. However, depending on the DSP cost, distance regimes where DV-QKD is better may appear, at it can be seen for the higher DSP costs. Even for the lowest value of DSP, BB84 with SNSPDs start to become more efficient after 387km.
@@
Finally at a last remark on Figure 7, note that this assumes that the noise parameters and reconciliation efficiency for CV-QKD are independent of the distance, which might not be true in practice. As an indication, the current distance records for CV-QKD are 202.81 km [33] using a shared local oscillator and square pulses, and 100 km [34, 35] using the newer pulse shaping methods, digital signal processing and local local oscillator. The main limitations come from imperfect phase noise correction, especially in the case of a local local oscillator, and limitations of error reconciliation at low signal-to-noise ratios. For DV-QKD, the distance record for a prepare-and-measure one-way protocol in an optical fiber is 421 km [36] using a three-state protocol with time-bin encoding.
IV Energy cost of multipartite protocols
Protocols involving three or more nodes in a quantum network are of particular interest for future deployments of the quantum internet. Not all quantum technologies scale linearly energy-wise, and therefore it is worth finding optimal configurations for multipartite networks. Firstly, different methods of generation of all-to-all entanglement are described and compared. Secondly, Conference Key Agreement (CKA) is used as a figure of merit for multipartite networks by comparing the performances of a few protocols constructed from DV and CV sources.
IV.1 All-to-all entanglement
@@
GHZ states are prime candidates for network applications since they allow the sharing of entanglement between all nodes at once [10, 12, 39]. A common photonic GHZ state creation setup involves SPDC sources creating Bell pairs that go through a fusion operation which is inherently probabilistic (see Appendix B.3). As a consequence, the probability of successfully creating a GHZ state of $n$ qubits decreases exponentially with $n$.
@@
$\begin{split}E_{\rm GHZ}(n,t)&=\left\lceil{\frac{n}{2}}\right\rceil\;S_{\textrm{SPDC}}(t)\\ &\hskip 11.38092pt+\left\lfloor{\frac{n-1}{2}}\right\rfloor\;M_{\rm fusion,polarization}(t)\\ &\hskip 17.07182pt+n\;D_{\rm SNSPD}(t)+C_{\rm GHZ}(t),\end{split}$ (7)
@@
$\begin{split}E_{\rm Alltoall}(n,t)&=\frac{n(n-1)}{2}\;S_{\textrm{SPDC}}(t)+n(n-1)\;M_{\rm polar}(t)\\ &\hskip 17.07182pt+n\;D_{\rm SNSPD}(t)+C_{\rm Alltoall}(t).\end{split}$ (8)
While this second method amounts to more hardware involved, the low probability of success of high-order GHZ state creation requires more repetitions and thus longer running times of the hardware. As shown in Figure 8, the best method to share entanglement between all nodes of a network varies with the number of parties. After $n=6$ parties, the probability of successfully creating a GHZ state becomes so low that it is less energetic demanding to use a pair-wise entangled architecture for this task.
@@
IV.2 Conference key agreement
Conference key agreement (CKA) [40] is the multipartite extension of QKD, allowing $n$ parties to create and share a common secret key. Similarly to QKD, several protocols achieve CKA, each with different pros and cons in terms of rate, security bound, and energy consumption.
@@ IV.2.1 DV-CKA
The most straightforward method to create a secret key between $n$ users of a quantum network is to distribute the qubits of a GHZ state to each party of a network. By measuring their qubits in the appropriate bases, the parties automatically extract a common private bit. A full security proof and description of a GHZ-based CKA protocol can be found in [41], while an experimental realization can be found in [42]. The rate of this protocol, which we denote GHZ-CKA, is given by the rate of creating, sharing, and measuring GHZ states that were distributed to the $n$ parties. Assuming that all $n$ parties are at an equal distance $d$ from the source of GHZ states, Equation 4 becomes:
@@ IV.2.2 CV-CKA
A brief description of the method to compute the secret key rate of the protocol [43] is given in Appendix C.2.3. As a concluding remark, each of the Bobs needs to perform post-processing of the measured outcomes, which means that the classical cost increases with the number of users (see Section III.3).
@@ IV.2.3 Simulation results for CKA
[…]
Figure 9: Energetic analysis of DV-CKA protocols: (a) Energy consumption of the three different CKA setups as a function of the number of parties for a fixed distance between the parties and the central node of 5 km, in log-scale. (b) Energy consumption of the three different CKA setups as a function of the distance between the parties and the central node, for a fixed number of parties $n=5$, in log-scale.
Figure 10: Energetic analysis of CV-CKA protocols. Main plot: comparison of the energy to distill 1Gbit of key with CV-CKA (plain) and nCV (dashed) protocols as a function of the distance to the central node (short distances only) for different number of users. Inset: Energy consumption required to distill 1Gbit of key with the nCV-QKD protocol (using homodyne measurements and double-polarization) as a function of the distance with the central node, for different numbers of users for larger distances.
In Figures 9 and 9, the energy cost $E^{\textrm{ \rm 1Gbit}}_{\textrm{CKA}}$ required to create 1 Gbit of key is illustrated, using the three aforementioned DV-CKA protocols as a function of the number of parties and as a function of the distance with the middle party Alice. The BB84-CKA protocol is always more energy efficient than the other two options. This is due to the high rate and high success probability of the BB84 protocol. As for the All-to-all task, there is a regime for both the number of parties and for the distance for which the GHZ-CKA protocol is more energy efficient than the Bell-CKA protocol. For longer distances and higher numbers of parties, the probability of a successful generation and detection of a GHZ state decreases to the point where it becomes more efficient to use bipartite communications between Alice and each Bob to accomplish this task.
@@
The energy consumption required to create 1 Gbit of secret key using the CV-CKA protocol of [43] as a function of distance is shown in Figure 10 for different numbers of parties. Regarding the comparison between continuous and discrete variables, the energy consumption of the CV-CKA protocol is three orders of magnitude lower than its DV-counterpart, but with major limitations on the achievable distance since this energetic advantage is valid only over a few hundred meters [43]. This is therefore a solution to be considered for a few users within building-scale distances, compared to the DV-CKA protocols presented in previous sections.
Figure 10 also shows the energy consumption required to create 1 Gbit of secret key using the nCV-QKD protocol for diverse numbers of users $n$, all of them separated by the same distance. A similar energy consumption with respect to the CV-CKA approach is observed with a noticeable improvement in terms of achievable distances (see inset). However, at the scale of few hundred meters, the CV-CKA approach performs better than nCV-QKD, hinting that such protocols could prove useful in access networks or buildings. Both techniques employ vastly different classical post-processing (in particular, the DSP), such that adding those contributions could modify the results of this simulation.
V Conclusion
In this article, we laid the foundations of a framework to estimate the energy consumption of quantum communication protocols. Two figures of merit were introduced, namely the energy efficiency and the energy cost required to produce a target number of secret bits. We applied them to different implementations of bipartite and multipartite protocols. The energy efficiency gives an idea of the consumption of a running network while the energy cost of producing a target number of bits gives a benchmark for the current and near-term energy cost of network protocols. By studying the energy cost required to solve the same task, hardware and protocol choices can be optimized.
This first insight into the energetic cost of photonic-based quantum communication protocols, including multipartite scenarios, shows that the critical components in discrete variable protocols are cryogenic-based hardware, while continuous variable protocols are deeply affected by post-processing. An interesting highlight of this study is that, for a distance of 25 km, a typical BB84 setup could generate 1 Gbit of secret key 60 times more efficiently using a less energy costly but less efficient detector than the usual cryogenic-based detectors, at the cost of 3 times the temporal requirements. These differences are critical at this early stage of quantum network development and can direct future efforts in different ways.